OpenStudy (anonymous):

The circumference of a circle is 50.24 centimeters. Find the area of the circle. (Use 3.14 for π.)

OpenStudy (anonymous):

7925.54 sq cm.

OpenStudy (angela210793):

@FoolForMath assuming ur right :)

OpenStudy (anonymous):

@angela210793: A fool can never be right :P

OpenStudy (anonymous):

Assuming it was rounded would u said it would be 803.84?

OpenStudy (anonymous):

No I don't think we can round to that value.

OpenStudy (anonymous):

Then your answer is wrong. The book has given me 4 answers. & That is not one of them. Please check your work, Explaining how u did it

OpenStudy (anonymous):

what is the area of the circle?

OpenStudy (anonymous):

πr2

OpenStudy (anonymous):

Would you like to see the answers that they gave to me?

OpenStudy (anonymous):

Somehow I got 200.845

OpenStudy (anonymous):

do you have 7930 as one option?

OpenStudy (anonymous):

These are the options the book gave to me. 50.24 cm2 803. 84 cm2 100.48 cm2 200.96 cm2

OpenStudy (anonymous):

GOD, this is circumference lol

OpenStudy (anonymous):

Blind me :(

OpenStudy (anonymous):

x] Its alright. Everyone makes mistakes.

OpenStudy (anonymous):

2 pi r = 50.24 r = 25.12/ pi area = pi r^2 area = pi (25.12)^2 / pi^2 area = (25.12)^2 / pi

OpenStudy (anonymous):

so it is 200.96 cm2

OpenStudy (anonymous):

circumference,C = 2 *pi*r = 50.24 pi*r=50.24/2 = 25.12 \[area = (Pi times r)^{2}\/Pi] area = 25.12*25.12/3.14= 200.96... This is the correct answer

OpenStudy (anonymous):

Thanks lol :]

OpenStudy (anonymous):

the above answers shocked me........ then i checked my answer twice.................